# Enabling Robust In-Context Memory and Rapid Task Adaptation in Transformers with Hebbian and Gradient-Based Plasticity

### v1

@@ 1 Introduction
Transformers exhibit striking in-context learning (ICL) capabilities: given a few input–output examples, they can perform novel tasks without weight updates. Yet this ability is largely an emergent consequence of scale and training diversity rather than an explicit mechanism for adaptation (Brown et al., 2020; Dai et al., 2023). As a result, the model’s capacity to incorporate new evidence is limited to transient activations within self-attention, with no dedicated process for storing or consolidating information during inference.
Biological neural systems, in contrast, continuously modify synaptic strengths in response to experience (Magee & Grienberger, 2020). Such synaptic plasticity enables rapid, context-dependent learning and flexible generalization (Bittner et al., 2017). Inspired by this principle, we ask: Can explicit plasticity mechanisms improve a Transformer’s ability to adapt within a sequence?
To answer this, we augment decoder-only Transformers with fast-weight components that are updated online at each generation step. These updates follow one of two biologically motivated rules: (i) a neuromodulated Hebbian rule, which adjusts weights locally based on correlated pre- and post-synaptic activity, and (ii) a gradient-based plasticity rule (Duan et al., 2023), which performs local gradient descent on an internally generated target signal. This formulation allows Transformers to adapt during inference, not just across training steps.
@@ 2.1 Meta-Learning for Rapid Adaptation
Our work falls under the "learning to learn" or meta-learning paradigm Schmidhuber et al. (1997); Bengio et al. (1992), which aims to build models that can adapt to new tasks from limited experience. This field is characterized by a "bi-level" optimization process: an "inner loop" where the model adapts to a specific task, and an "outer loop" where the meta-learner is updated to improve this adaptation process. Several distinct approaches have emerged:
  • Gradient-based Meta-Learning: This line of work, exemplified by Model-Agnostic Meta-Learning (MAML) Finn et al. (2017), meta-learns a parameter initialization $\theta$ that can be effectively fine-tuned to a new task with only a few gradient steps. This approach is powerful but typically assumes that the inner-loop task provides an explicit, differentiable loss function.
  • Metric-based Meta-Learning: Methods like Prototypical Networks Snell et al. (2017) learn a deep embedding space where new examples can be classified based on their proximity to "prototypes" (e.g., class means) derived from the support set. While effective, these methods are often specialized for few-shot classification.
  • Memory-Augmented Networks (MANNs): These models, such as those proposed by Santoro et al. Santoro et al. (2016), utilize an external memory (e.g., a Neural Turing Machine’s memory bank) to store task-specific information, which the controller network can read from and write to.
  • Meta-Learning a Learning Rule: Our work, following Duan et al. Duan et al. (2023), belongs to a category that meta-learns the learning rule itself. Instead of learning an initialization (like MAML) or an embedding (like ProtoNets), the outer loop meta-trains the parameters of a dynamic update rule (e.g., the connection-specific learning rates $\alpha$).
@@ 2.2 Dynamic Parameters and Synaptic Plasticity
  • Biologically-Inspired Plasticity: This approach draws direct inspiration from neuroscience. Hebb’s rule ("cells that fire together, wire together") is a classic model of such plasticity. In ANNs, this has been adapted to create associative memories Limbacher & Legenstein (2020); Schlag et al. (2021b) and to meta-learn policies in simple reinforcement learning tasks Najarro & Risi (2020). The most direct precursors to our work are Miconi et al. Miconi et al. (2018; 2019) and the original paper by Duan et al. Duan et al. (2023), which applied differentiable, generalized Hebbian rules to RNNs to solve memory and meta-learning problems.
  • Hypernetworks and Fast Weight Programmers: An alternative, engineering-driven approach is to use a secondary network, or "hypernetwork," to generate the weights of a primary network Ha et al. (2017). This concept has been adapted for Transformers, notably in "Fast Weight Programmers" Schlag et al. (2021a); Irie et al. (2021), where a recurrent network or attention mechanism generates the fast weights used by another component.
Our work differs from this latter branch by not using a separate network to generate weights. Instead, we implement specific, local (Hebbian) and global (gradient-based) update rules that modify the weights based on the network’s own activity, as proposed in Duan et al. (2023).
@@
A significant challenge and inspiration for this work is the remarkable in-context learning (ICL) ability of large-scale Transformers Brown et al. (2020). These models appear to learn new tasks at inference time from examples in their context, all without any explicit fast weights or gradient updates. This has led to two diverging views on the mechanism of ICL:
  • The "Implicit" View (ICL as Optimization): A prominent theoretical line analyzes the Transformer’s forward pass as an implicit optimization algorithm. From this perspective, the self-attention mechanism learns to implement a learning algorithm (e.g., a form of gradient descent Garg et al. (2022) or Bayesian inference Xie et al. (2022)) on the context. The model’s parameters (slow weights) are trained to encode an algorithm, such as ridge regression, which is then executed in the forward pass using the context examples Akyürek et al. (2023). In this view, "learning" is an ephemeral process that exists only in the activation patterns and attention matrices.
  • The "Explicit" View (ICL as Plasticity): This is the "dense-versus-sparse" taxonomy we explore. The implicit view describes a "sparse" update, where learning is encoded in activations. We investigate the "dense" alternative: modifying the network’s weights themselves. Our work tests whether augmenting a Transformer with the explicit plasticity mechanisms from Duan et al. Duan et al. (2023) provides a more direct, robust, or efficient mechanism for ICL.
@@ 3.2 Model Framework
Plastic Transformers are trained using a two-timescale optimization loop (Algorithm 1). Within each sequence (the inner loop), the fast weights are updated at every step using either a Hebbian or gradient-based plasticity rule. Across sequences, an outer loop optimizes the static parameters $\tilde{W}_{l}$ and connection-specific learning rates $\alpha_{l}$ to minimize the overall meta-loss. This setup parallels learned-optimizer frameworks and the plastic-RNN training of Duan et al. (2023).
@@
At each step, the model outputs both the token prediction $y_{t}$ and auxiliary variables $(\tilde{\eta}_{t},\overline{y}_{t})$ that control the internal modulation and gradient-based updates. The additional memory required for storing plastic weights scales as $O(d_{\text{model}}^{2})$ per plastic layer, but in practice contributes less than 5% computational overhead relative to a non-plastic baseline.
@@ 3.3 Hebbian Plasticity
where $\alpha_{l}$ are learnable per-connection learning rates and $\eta(t)$ is a global, learned neuromodulation factor controlling when plasticity is active. Following Duan et al. (2023), we compute
@@ 3.4 Gradient-Based Plasticity
Here $\beta_{l}$ are learnable bias-specific rates, and $\eta(t)$ is computed as in Eq. 3 with $\delta_{t}$ defined as the concatenation of all internal gradients. This rule gives the model an explicit mechanism for credit assignment within the current sequence, complementing the associative memory provided by the Hebbian rule.
@@ 4 Experiments
We evaluate plastic Transformers on the same suite of tasks introduced by Duan et al. (2023), testing two hypotheses: (1) explicit plasticity improves short-term memory, and (2) it enables rapid, in-sequence learning. All data generation, losses, and evaluation protocols follow the original setups, with model sizes matched in parameter count to ensure comparability.
@@ 4.1 Experimental Protocol
Each configuration is run with $S=3$ independent seeds. Reported values are mean $\pm$ standard deviation across seeds. Every runner includes a non-plastic baseline (rule=none) trained with identical hyperparameters, isolating the contribution of the fast-weight mechanism. Dataset sizes are tuned for computational efficiency: 50 sequences (copying), 100 cue–reward trials, 150 regression episodes, and 80/100 classification episodes (CIFAR-FS/Omniglot). Doubling dataset sizes in pilot sweeps did not alter the qualitative ranking of rules.
@@ 4.4 One-Shot Image Classification
Task. 5-way, 1-shot sequential classification on CIFAR-FS and Omniglot. Setup. A Conv-4 encoder provides embeddings consumed sequentially by the Transformer. Metric. Validation accuracy (higher is better).
@@
Interpretation. Hebbian plasticity achieves the best accuracy on both datasets (CIFAR-FS: +4.6 pp; Omniglot: +3.6 pp over gradient). Gradient updates tend to maintain large, persistent $\eta(t)$ values ($0.116$ / $0.061$) that degrade accuracy by overfitting to sparse supervision, while Hebbian plasticity fires only around labelled supports. This aligns with the hypothesis that local, associative updates outperform dense gradient surrogates in sparse, class-conditional regimes. For reference, plastic RNNs from Duan et al. (2023) achieve 55.5 ± 1.0
@@ 4.6 Comparison with Plastic RNNs
To contextualize Transformer results, Table 5 compares performance against plastic RNNs from Duan et al. (2023) on identical benchmarks.
Table 5: Cross-architecture comparison with plastic RNNs (Duan et al., 2023).
@@ 5 Discussion
The observed task-dependent behaviour aligns with theoretical accounts of in-context optimization in Transformers (Akyürek et al., 2023; von Oswald et al., 2023). Dense credit assignment regimes favour global, gradient-like update mechanisms, whereas sparse or discrete supervision naturally engages local correlation-based rules. In this sense, the two plasticity forms occupy complementary positions along a continuum from implicit in-context learning to explicit fast-weight adaptation. The results also reinforce that explicit plasticity can coexist stably with self-attention and provides a controllable handle on when and how adaptation occurs within a sequence.
@@ 6 Code and Compute
All experiments were implemented in Python 3.10 using PyTorch 2.1.1 and the Hydra configuration framework. Each benchmark (copying, cue–reward, few-shot regression, CIFAR-FS, and Omniglot classification) is defined as an independent configuration module, with runs launched via unified command-line wrappers across three random seeds (3000–3002). Raw training logs, per-epoch diagnostics, and aggregated metrics are automatically stored as JSON files, enabling full reproducibility. All figures and tables were generated by a single plotting script that consumes these logs and outputs publication-ready assets. The complete codebase, including configurations and pretrained checkpoints, is available at the project repository: https://github.com/sidcraftscode/hebbian-transformer.
@@ Appendix B Code Availability
https://github.com/sidcraftscode/hebbian-transformer

### v2

@@ 1 Introduction
Transformers exhibit striking in-context learning (ICL) capabilities: given a few input–output examples, they can perform novel tasks without weight updates. Yet this ability is largely an emergent consequence of scale and training diversity rather than an explicit mechanism for adaptation (brown2020language; dai2023can). As a result, the model’s capacity to incorporate new evidence is limited to transient activations within self-attention, with no dedicated process for storing or consolidating information during inference.
Biological neural systems, in contrast, continuously modify synaptic strengths in response to experience (magee2020synaptic). Such synaptic plasticity enables rapid, context-dependent learning and flexible generalization (bittner2017behavioral). Inspired by this principle, we ask: Can explicit plasticity mechanisms improve a Transformer’s ability to adapt within a sequence?
To answer this, we augment decoder-only Transformers with fast-weight components that are updated online at each generation step. These updates follow one of two biologically motivated rules: (i) a neuromodulated Hebbian rule, which adjusts weights locally based on correlated pre- and post-synaptic activity, and (ii) a gradient-based plasticity rule (duan2023hebbian), which performs local gradient descent on an internally generated target signal. This formulation allows Transformers to adapt during inference, not just across training steps.
@@ 2.1 Meta-Learning for Rapid Adaptation
Our work falls under the "learning to learn" or meta-learning paradigm schmidhuber1997shifting; bengio1992optimization, which aims to build models that can adapt to new tasks from limited experience. This field is characterized by a "bi-level" optimization process: an "inner loop" where the model adapts to a specific task, and an "outer loop" where the meta-learner is updated to improve this adaptation process. Several distinct approaches have emerged:
  • Gradient-based Meta-Learning: This line of work, exemplified by Model-Agnostic Meta-Learning (MAML) finn2017model, meta-learns a parameter initialization $\theta$ that can be effectively fine-tuned to a new task with only a few gradient steps. This approach is powerful but typically assumes that the inner-loop task provides an explicit, differentiable loss function.
  • Metric-based Meta-Learning: Methods like Prototypical Networks snell2017prototypical learn a deep embedding space where new examples can be classified based on their proximity to "prototypes" (e.g., class means) derived from the support set. While effective, these methods are often specialized for few-shot classification.
  • Memory-Augmented Networks (MANNs): These models, such as those proposed by Santoro et al. santoro2016meta, utilize an external memory (e.g., a Neural Turing Machine’s memory bank) to store task-specific information, which the controller network can read from and write to.
  • Meta-Learning a Learning Rule: Our work, following Duan et al. duan2023hebbian, belongs to a category that meta-learns the learning rule itself. Instead of learning an initialization (like MAML) or an embedding (like ProtoNets), the outer loop meta-trains the parameters of a dynamic update rule (e.g., the connection-specific learning rates $\alpha$).
@@ 2.2 Dynamic Parameters and Synaptic Plasticity
  • Biologically-Inspired Plasticity: This approach draws direct inspiration from neuroscience. Hebb’s rule ("cells that fire together, wire together") is a classic model of such plasticity. In ANNs, this has been adapted to create associative memories limbacher2020harnessing; schlag2021b and to meta-learn policies in simple reinforcement learning tasks najarro2020meta. The most direct precursors to our work are Miconi et al. miconi2018differentiable; miconi2019backpropamine and the original paper by Duan et al. duan2023hebbian, which applied differentiable, generalized Hebbian rules to RNNs to solve memory and meta-learning problems.
  • Hypernetworks and Fast Weight Programmers: An alternative, engineering-driven approach is to use a secondary network, or "hypernetwork," to generate the weights of a primary network ha2017hypernetworks. This concept has been adapted for Transformers, notably in "Fast Weight Programmers" schlag2021linear; irie2021going, where a recurrent network or attention mechanism generates the fast weights used by another component.
Our work differs from this latter branch by not using a separate network to generate weights. Instead, we implement specific, local (Hebbian) and global (gradient-based) update rules that modify the weights based on the network’s own activity, as proposed in duan2023hebbian.
Relation to test-time training (TTT).
Our gradient-based plasticity is related in spirit to test-time adaptation methods that update parameters using an internally defined objective during inference (often called test-time training or adaptation). However, there are key differences: (i) our inner objective $L(t)$ is computed per step from variables already emitted by the model and targets only the fast plastic state; (ii) the outer objective remains the task loss and trains the rules and slow parameters that make such step-wise adaptation useful; and (iii) our updates are neuromodulated and gated, not unconditional. Recent work on TTT (dalal2025tttvideo; behrouz2024titans) typically optimizes a self-supervised auxiliary loss on the same slow parameters at test time, whereas here only fast weights are updated online while slow weights remain fixed.
@@
A significant challenge and inspiration for this work is the remarkable in-context learning (ICL) ability of large-scale Transformers brown2020language. These models appear to learn new tasks at inference time from examples in their context, all without any explicit fast weights or gradient updates. This has led to two diverging views on the mechanism of ICL:
  • The "Implicit" View (ICL as Optimization): A prominent theoretical line analyzes the Transformer’s forward pass as an implicit optimization algorithm. From this perspective, the self-attention mechanism learns to implement a learning algorithm (e.g., a form of gradient descent garg2022transformers or Bayesian inference xie2022explanation) on the context. The model’s parameters (slow weights) are trained to encode an algorithm, such as ridge regression, which is then executed in the forward pass using the context examples akyurek2023learning. In this view, "learning" is an ephemeral process that exists only in the activation patterns and attention matrices.
  • The "Explicit" View (ICL as Plasticity): This is the "dense-versus-sparse" taxonomy we explore. The implicit view describes a "sparse" update, where learning is encoded in activations. We investigate the "dense" alternative: modifying the network’s weights themselves. Our work tests whether augmenting a Transformer with the explicit plasticity mechanisms from Duan et al. duan2023hebbian provides a more direct, robust, or efficient mechanism for ICL. Note that standard Transformers use only the implicit mechanism; the explicit variant refers to our plastic-augmented model.
@@ 3.2 Model Framework
Plastic Transformers are trained using a two-timescale optimization loop (Algorithm 1). Within each sequence (the inner loop), the fast weights are updated at every step using either a Hebbian or gradient-based plasticity rule. Across sequences, an outer loop optimizes the static parameters $\tilde{W}_{l}$ and connection-specific learning rates $\alpha_{l}$ to minimize the overall meta-loss. This setup parallels learned-optimizer frameworks and the plastic-RNN training of duan2023hebbian.
@@
At each step, the model outputs both the token prediction $y_{t}$ and auxiliary variables $(\tilde{\eta}_{t},\overline{y}_{t})$ that control the internal modulation and gradient-based updates. Here, $\tilde{\eta}_{t}$ is a scalar neuromodulation logit (pre-sigmoid) and $\overline{y}_{t}\in\mathbb{R}^{d_{\text{aux}}}$ is an auxiliary vector produced by a small head alongside $y_{t}$. The additional state for plasticity consists of the fast weights $w_{l}(t)$ (and biases) matching the shape of the corresponding static weights, i.e., $O(d_{\text{model}}^{2})$ parameters per plastic FFN matrix. This adds a small, task- and configuration-dependent overhead in runtime and memory; the code paths make this explicit, and we quantify sizes below when discussing each rule.
@@ 3.3 Hebbian Plasticity
where $\alpha_{l}$ are learnable per-connection learning rates and $\eta(t)$ is a global, learned neuromodulation factor controlling when plasticity is active. The operator $\circ$ denotes elementwise (Hadamard) product and $\text{Vec}(\cdot)$ flattens a matrix. Following duan2023hebbian, we compute
@@ 3.4 Gradient-Based Plasticity
Here $\beta_{l}$ are learnable bias-specific rates, and $\eta(t)$ is computed as in Eq. 3 with $\delta_{t}$ defined as the concatenation of all internal gradients. $W_{\text{out}}\in\mathbb{R}^{d_{o}\times d_{o}}$ is a learned square projection mapping the concatenated vector $\text{Concat}(y_{t},\overline{y}_{t},\tilde{\eta}_{t})\in\mathbb{R}^{d_{o}}$ back to $\mathbb{R}^{d_{o}}$; here $d_{o}=d_{\text{class}}+d_{\text{aux}}+1$. This rule gives the model an explicit mechanism for credit assignment within the current sequence, complementing the associative memory provided by the Hebbian rule.
Computational complexity and overhead.
Memory overhead comes from storing fast weights $w_{l}(t)$ (and biases) with the same shape as their static counterparts. For a plastic FFN with two linear maps of shapes $(d_{\text{ff}},d_{\text{model}})$ and $(d_{\text{model}},d_{\text{ff}})$, the additional state per block is $O(d_{\text{ff}}\,d_{\text{model}})$ parameters per matrix (plus optional biases). Hebbian updates add outer-products and elementwise operations per step (roughly $O(d_{\text{ff}}\,d_{\text{model}})$ per plastic matrix). The gradient-based rule additionally computes an internal projection on $d_{o}$-dimensional vectors (cost $O(d_{o}^{2})$ per step, typically small) and an autograd pass to obtain $\partial L/\partial w$, whose cost is comparable to a lightweight backward through the plastic FFN. In all cases, slow (static) parameters are unchanged during inference; only the fast state is updated online.
@@ 4 Experiments
We evaluate plastic Transformers on the same suite of tasks introduced by duan2023hebbian, testing two hypotheses: (1) explicit plasticity improves short-term memory, and (2) it enables rapid, in-sequence learning. All data generation, losses, and evaluation protocols follow the original setups, with model sizes matched in parameter count to ensure comparability.
@@ 4.1 Experimental Protocol
Each configuration is run with S=3S{=}3 independent seeds. Reported values are mean $\pm$ standard deviation across seeds.
Evaluation isolates plasticity: every experiment includes a non-plastic baseline (rule=none) trained with identical optimizer, architecture, and data protocol.
Inner vs. outer updates. Within each sequence/episode (inner loop), the plastic state is initialized to zero and updated online at every step via either the Hebbian increment or the gradient-based step on $L(t)$. The static parameters (including $\tilde{W}$ and the learning-rate tensors $\alpha,\beta$) are updated only by the outer loop, once per sequence, using the task loss accumulated across that sequence. At test time, no outer-loop updates are applied: models adapt only through their in-sequence plasticity.
Data generation and dataset sizes. For synthetic tasks (copying, cue–reward, regression), episodes are generated on the fly in __getitem__, so there is no fixed set to memorize; dataset_size controls the number of fresh sequences drawn per epoch. For the reported runs we use: 50 sequences (copying; delay 20), 100 cue–reward trials, 150 regression episodes, and 80/100 classification episodes per epoch (CIFAR-FS/Omniglot). Each sequence is seen once per epoch for $E$ epochs (see schedule below), with state reset between sequences. This balances compute and variance while keeping protocols consistent across rules.
Inner update counts per task. The number of inner (plastic) updates per episode equals the sequence length $T$:
  • Copying: $T=2\,\text{seq\_length}+\text{delay}+1$ (presentation, delay, delimiter, recall).
  • Cue–reward: $T=2\,\text{num\_pairs}$ (present then query for each cue).
  • Few-shot regression: $T=K_{\text{support}}+K_{\text{query}}$.
  • Few-shot classification: $T=N\!\times\!K$ (support) $N\!\times\!Q$ (queries) for $N$-way, $K$-shot, $Q$-query.
Per-task training schedule (as used in released artifacts). Copying: 50 episodes/epoch for E=2E{=}2 epochs (total $100$ episodes/run). Cue–reward: 100 episodes/epoch, E=4E{=}4. Regression: 150 episodes/epoch, E=5E{=}5. CIFAR-FS: 80 episodes/epoch, E=5E{=}5. Omniglot: 100 episodes/epoch, E=5E{=}5. Outer-loop updates occur once per episode; inner updates occur $T$ times per episode as above.
@@ 4.4 One-Shot Image Classification
[…]
@@
Interpretation. Hebbian plasticity achieves the best accuracy on both datasets (CIFAR-FS: +4.6 pp; Omniglot: +3.6 pp over gradient). Gradient updates tend to maintain large, persistent $\eta(t)$ values ($0.116$ / $0.061$) that degrade accuracy by overfitting to sparse supervision, while Hebbian plasticity fires only around labelled supports. This aligns with the hypothesis that local, associative updates outperform dense gradient surrogates in sparse, class-conditional regimes. For reference, plastic RNNs from duan2023hebbian achieve 55.5 ± 1.0
@@ 4.6 Comparison with Plastic RNNs
To contextualize Transformer results, Table 5 compares performance against plastic RNNs from duan2023hebbian on identical benchmarks.
Table 5: Cross-architecture comparison with plastic RNNs (duan2023hebbian).
@@ 5 Discussion
The observed task-dependent behaviour aligns with theoretical accounts of in-context optimization in Transformers (akyurek2023learning; vonoswald2023transformers). Dense credit assignment regimes favour global, gradient-like update mechanisms, whereas sparse or discrete supervision naturally engages local correlation-based rules. In this sense, the two plasticity forms occupy complementary positions along a continuum from implicit in-context learning to explicit fast-weight adaptation. The results also reinforce that explicit plasticity can coexist stably with self-attention and provides a controllable handle on when and how adaptation occurs within a sequence.
@@ 6 Code and Compute
All experiments were implemented in Python 3.10 using PyTorch 2.1.1 and the Hydra configuration framework. Each benchmark (copying, cue–reward, few-shot regression, CIFAR-FS, and Omniglot classification) is defined as an independent configuration module, with runs launched via unified command-line wrappers across three random seeds (3000–3002). Raw training logs, per-epoch diagnostics, and aggregated metrics are automatically stored as JSON files, enabling full reproducibility. All figures and tables were generated by a single plotting script that consumes these logs and outputs publication-ready assets. The complete codebase, including configurations and pretrained checkpoints, is available at the project repository: https://github.com/sidcraftscode/hebbian-transformer/.
@@ Appendix B Code Availability
https://github.com/sidcraftscode/hebbian-transformer/